OpenStudy (minecraft40952):

you deposit $250.00 in an account that earns 6% compounded annually (once per year). What is the Balance in your account after 3 years? Round your Answer to the nearest cent. I just need the equation. i cant get it in equation form

OpenStudy (minecraft40952):

@confluxepic Pls help

OpenStudy (minecraft40952):

@dan815 pls help

OpenStudy (anonymous):

$250.00/ 18%. i think.

OpenStudy (minecraft40952):

no i need it in equation form not division

OpenStudy (minecraft40952):

like y=mx+b

OpenStudy (minecraft40952):

stuff like that

OpenStudy (anonymous):

m=a-p?? i know the answer just don't quite get the equation part

OpenStudy (minecraft40952):

Wat M=a-p im talking about Y=Mx+b

OpenStudy (anonymous):

aye it was a idea.... sorry

OpenStudy (minecraft40952):

no ive had a bad day no one helps anymore they sit in the chat rooms and talk about personal life and drugs and crap

OpenStudy (minecraft40952):

While people like me need help

OpenStudy (anonymous):

yes i understand... people do help, you just need to find the right people. and you need it in equation form? so like 205=250-18%

OpenStudy (minecraft40952):

oh i just figured it out heres a medal and fan for actruly trying to help! :D

OpenStudy (anonymous):

OK thanks